OpenStudy (anonymous):

Please help! f (x) = 1/3 x + 3 g (x) = x - 2 The functions f and g above are defined for all values of x. In the xy-plane, what is the y-coordinate of the point of intersection of the graphs of f and g?

OpenStudy (anonymous):

Well, where do f and g intersect? Set them equal to each other: (1/3)x + 3 = x - 2 (2/3)x = 5 x = 7.5 So if x = 7.5, the y-coordinate is found by substituting x = 7.5 into any one of the given euations, say f. you get (1/3)(7.5) + 3 = 5.5